# Learning to Construct Implicit Communication Channel

### v1

@@ 6 Experiments
In this section, we explore the application of the ICP across various environments where no direct communication is available. ICP plays a vital role in these settings by constructing an implicit communication channel that enhances the agents’ ability to share and interpret information. We will demonstrate how ICP facilitates effective collaboration and decision-making in 3 tasks, namely Guessing Number, the Revealing Goals, and the Hanabi card game. The first two tasks are designed by us and can be reused by future works as testing environments. The Hanabi game is a popular card game played by humans, and is described as The Hanabi Challenge (Bard et al., 2020) by the community. Each of these environments presents unique challenges, and we will show how ICP optimizes the use of scouting actions to improve overall performance and strategy.
Figure 1: Left: Guessing Number Environment . Agents can’t see their own digits, but can reveal others’ segments’ state, by collaboratively give hints, they can deduce their own digit and obtain shared rewards. Right: Revealing Goals Environment. Each agents are positioned a random position in a grid world and are assigned a unique target. However, they can only observe others’ targets which not include themselves. By revealing each other’s targets in the nearby grid, they can eventually find their own targets and reach them.
6.1 Guessing Number
In the Guessing Number Game, we present a collaborative, turn-based multi-agent reinforcement learning environment involving $N$ agents. Each agent $i$ is uniquely assigned a digit $d_{i}\in\{0,1,\dots,9\}$, which is visible to all other agents $j\neq i$ but remains unknown to agent $i$ themselves. The primary objective for each agent is to deduce and correctly guess their own digit $d_{i}$. During each turn, an agent can choose to either guess their own digit by selecting $\hat{d}_{i}\in\{0,1,\dots,9\}$ or provide a hint about another agent’s digit using the Radioland Slim font—a seven-segment digital display representation. Specifically, hints involve revealing the state (“on” or “off”) of one of the seven segments of another agent’s digit, and these hints are public and observable by all agents. Each agent’s action space comprises $10+7\times(N-1)$ actions: 10 options for guessing their own digit and $7$ hint options for each of the $N-1$ other agents. The game encourages strategic collaboration, as agents must balance between gathering information to deduce their own digit and assisting others through informative hints, aiming for the collective success of all agents correctly guessing their digits.
@@ 6.1 Guessing Number
In the experiment of Guessing Number, we compare the performance of VDN-on-policy and VDN-off-policy with ICN (Implicit Channel Net, our implementation of ICP with a randomly initialized information map). Each approach is evaluated over 1k episodes and averaged over 6 random seeds. Each seed is running on a Linux metal machine with 256 GB RAM and 3090Ti GPU for 36 hours.
For VDN-off-policy, we begin by warming up the replay buffer until its size exceeds the batch size. During each training step, we add 10 episodes to the replay buffer and randomly sample batch-size episodes from the buffer for training. In contrast, for VDN-on-policy and our proposed method, we utilize a vectorized environment to sample batch size episodes at each training step and use these samples for training.
@@
The results is presented in Figure 2. All algorithms are capable of eventually guessing their own digit correctly, but the differences of number of hit actions needed is significant among the algorithms. By examining the saved model parameters and averaging over 10k episodes, we found that ICN has an average episode length of approximately 8.63, while VDN-on-policy and VDN-off-policy have average episode lengths of around 12.4. This observation further demonstrates that ICN achieves more efficient information transfer by using scouting actions to construct implicit channels.
Figure 2: (a): The training curves of Guessing Number over total 100k train steps with $N=3,l=11$. (b): Average episode length running by different algorithms in Guessing Number.
@@ 6.2 Revealing Goals
In the experiments for the Revealing Goals task, we used the same training hyperparameters and model architecture as in the Guessing Number experiments. We compared the performance of ICN with VDN-on-policy and VDN-off-policy methods. The experimental result is shown in Figure 3(a). ICN significantly outperforms the baseline methods by achieving an average score 2.17 times of the baselines. One reason for its effectiveness is that, in each round, each of the $N$ agents can engage in $N$ broadcast communications through implicit channels, which significantly enhances information efficiency compared to relying solely on environmental feedback. This result further validates the effectiveness of ICP in more complicated grid world tasks.
@@ 6.3 Hanabi
Because the game rule of Hanabi is quite involved, we refer it to a popular website https://hanabi.github.io/ that discusses both rules and human-play strategies. The action space for each agent includes three options: playing a card out of their 5-card hand, discarding a card, or giving a hint towards another player’s 5-card hand. The hint could reveal all cards in a player’s hand of a specific color or rank. Since they cannot see their own cards, agents must rely on hints from others to infer what cards they hold. The success of the game heavily depends on the agents’ ability to use these hints effectively to coordinate their actions and play the correct cards in order.
In our experiments, we focused on the 4-player full version of Hanabi. Since hint actions in Hanabi are natural scouting actions, we selected them to serve as information carriers for our ICP implementation. We implemented several baseline algorithms, including VDN-off-policy and the original version of SAD (Hu & Foerster, 2019), as well as a variant of SAD with the hint action masks removed. For the ICP algorithm, we explored several variations, including ICN-DIAL-on-policy, which involves strategy training with a random initial map using DIAL’s communication gradient technique resulting in an on-policy learning method. We also implemented ICN-DIAL-cut-gradient-off-policy, similar to ICN-DIAL-on-policy but with the communication gradient chain cut off to enable off-policy learning. Additionally, we used ICN-RGMComm, which is strategy training with a delayed map using RGMComm (Chen et al., 2024), an off-policy MARL communication algorithm, and ICN-RGMComm-hat-mapping, an extension of ICN-RGMComm that utilizes the hat mapping method for improved communication.
Our presented results are based on the best runs selected from more than three random seeds. From the results shown in Figure 3(b), we observe that the ICP method using the hat mapping approach, ICN-RGMComm-hat-mapping, achieved the best performance. The score achieved by ICP is 24.91, which surpasses the best available learning algorithm (SAD, 23.81 points) and vastly outperforms theory of mind-based methods (Fuchs et al. (2021) 19.13 points) and average human players (17 points, reported in Kantack (2021)). It is worth remarking that the state-of-the-art strategy in the Hanabi game is a search-based algorithm, WTFWThat+search, which reports an averaged score of $24.96$ points (Lerer et al., 2020). With the maximum score of the game being $25$, and there are certain deals that prevent the players to win $25$ points in any way, both ICP and WTFWThat+search are close to solving the game.
We also discussed the comparison between ICP and Hanabi human convention in Appendix B.
7 Discussion
@@ 7.1 Compatibility of ICP with Other Training Methods
In the ICP framework, action policy allows for the use of various algorithms for training. For instance, the action policy can be trained using value-based methods like VDN (Sunehag et al., 2018) or policy-based methods such as MADDPG (Lowe et al., 2017). This compatibility ensures that ICP can be tailored to various problem settings and agent dynamics, allowing researchers to choose the most suitable algorithm depending on the specific characteristics of the environment and the agents involved.
The communication strategy within the ICP framework is also versatile. It can be implemented using a variety of direct channel communication algorithms that support discrete channels. These include well-established methods like DIAL or RGMComm, which are designed to optimize communication under different constraints and requirements. The flexibility in choosing both the action policy and the communication algorithm highlights the comparability and broad applicability of the ICP framework, making it a general approach for multi-agent reinforcement learning scenarios.
@@ 7.2 Potential of Further Utilizing the Environment Information
Figure 4: By shuffling the embedding of information into scouting action, even if the information strategy stays the same, environment information will also change. After fine-tuning, the performance of implementation with the same information strategy but different embedding varies.
As described in Section 4, we analyzed that scouting actions carry two types of information. The ICP framework utilizes the information reflected through the choice of the scouting action (choice information) to establish an implicit communication channel. However, even when we leverage this information, the information reflected through the environment (environment information) does not disappear. Although in most cases the constructed implicit channel transmits more stable and useful information, in some situations the environment information is also significant. For example, in the Revealing Goals environment, a revealing action intended to transmit embedded information might also reveal other agents’ goals, which helps agents to make better decisions.
To verify the impact of this environmental information on performance, we randomly shuffled 6 different embeddings on the ICN-RGMComm-hat-mapping policy saved from the previous Hanabi experiment. We then fine-tuned the action policy while keeping the information strategy fixed. The results are shown in Figure 4, where, under a fixed information strategy, some randomly shuffled embeddings enjoy improved performance. This suggests that one could further utilize the environment information for more effective communication. We leave this research problem for future works.
@@ 8 Conclusion and Future Work
The code and our designed environments is freely available in the supplementary material. The details of the implementation can be found in Appendix A. The hyper-parameters used in experiments also mentioned in Section 6.
@@ Appendix B Human Convention in Hanabi
Compare with H-Group conventions, our ICP framework differs in that the communication strategy itself is learned as part of the training process, rather than being predefined. The communication through informative actions becomes akin to following a learned convention where agents develop their own patterns for encoding and transmitting information during the training phase. This allows agents to use scouting or informative actions as explicit communication mechanisms, while regular actions are selected through an online decision-making policy during gameplay. Notably, our method achieves a score of approximately $24.91$ in the Hanabi game, which significantly surpasses the average human score of around $17$ (Kantack, 2021). Thus, ICP offers a dynamic communication framework where the convention of communication emerges from the agents’ learning process, enabling more flexible and adaptive strategies depending on the environment and task at hand.

### v2

@@ 6 Experiments
In this section, we explore the application of the ICP across various environments where no direct communication is available. ICP plays a vital role in these settings by constructing an implicit communication channel that enhances the agents’ ability to share and interpret information. We will demonstrate how ICP facilitates effective collaboration and decision-making in 3 tasks, namely Guessing Numbers, the Revealing Goals, and the Hanabi card game. The first two tasks are designed by us and can be reused by future works as testing environments. The Hanabi game is a popular card game played by humans, and is described as The Hanabi Challenge (Bard et al., 2020) by the community. Each of these environments presents unique challenges, and we will show how ICP optimizes the use of scouting actions to improve overall performance and strategy.
Figure 1: Left: Guessing Numbers Environment . Agents cannot see their own digits, but can reveal others’ segments’ states by collaboratively giving hints. They can deduce their own digit and obtain shared rewards. Right: Revealing Goals Environment. The agents are positioned in a random spots in a grid world and are assigned a unique target. However, they can only observe others’ targets which do not include themselves. By revealing each other’s targets in the nearby grid, they can eventually find their own targets and reach them.
6.1 Guessing Numbers
In the Guessing Numbers game, we present a collaborative, turn-based multi-agent reinforcement learning environment involving $N$ agents. Each agent $i$ is uniquely assigned a digit $d_{i}\in\{0,1,\dots,9\}$, which is visible to all other agents $j\neq i$ but remains unknown to agent $i$ themselves. The primary objective for each agent is to deduce and correctly guess their own digit $d_{i}$. During each turn, an agent can choose to either guess their own digit by selecting $\hat{d}_{i}\in\{0,1,\dots,9\}$ or provide a hint about another agent’s digit using the Radioland Slim font—a seven-segment digital display representation. Specifically, hints involve revealing the state (“on” or “off”) of one of the seven segments of another agent’s digit, and these hints are public and observable by all agents. Each agent’s action space comprises $10+7\times(N-1)$ actions: 10 options for guessing their own digit and $7$ hint options for each of the $N-1$ other agents. The game encourages strategic collaboration, as agents must balance between gathering information to deduce their own digit and assisting others through informative hints, aiming for the collective success of all agents correctly guessing their digits.
@@ 6.1 Guessing Numbers
In the Guessing Numbers experiment, we evaluate the performance of 5 approaches: VDN-on-policy, VDN-off-policy, ICP with the random initial map approach (ICN-DIAL-RM), ICP with the delayed map approach (ICN-DIAL-DM), and a cheating approach where a direct communication channel is available (DIAL-Cheat). Each approach is evaluated over 1k episodes with 6 random seeds, and running on a Linux metal machine with 256 GB RAM and 3090Ti GPU for 36 hours.
For VDN-off-policy, we begin by warming up the replay buffer until its size exceeds the batch size. During each training step, we add 10 episodes to the replay buffer and randomly sample a batch of episodes from the buffer for training. In contrast, for VDN-on-policy and our proposed method, we utilize a vectorized environment to sample a batch of episodes at each training step and use these samples for training.
@@
The results are presented in Figure 2. All algorithms are capable of eventually guessing their own digit correctly, but the differences of number of hit actions needed is significant among the algorithms. By examining the saved models and averaging over 10k episodes, we find that ICN-DIAL-RM has an average episode length of approximately 8.63, while VDN-on-policy and VDN-off-policy have average episode lengths of around 12.4. This observation further demonstrates that ICN achieves more efficient information transfer by using scouting actions in implicit channels.
Figure 2: (a): The training curves of Guessing Numbers over a total of 100k steps with $N=3,l=11$. (b): Average episode length running by different algorithms in Guessing Numbers.
@@ 6.2 Revealing Goals
In the experiments for the Revealing Goals task, we use the same training hyperparameters and model architecture as in the Guessing Numbers experiments. We also compare the performance of 5 approaches: DIAL-Cheat, ICN-DIAL-DM, ICN-DIAL-RM, VDN-on-policy, and VDN-off-policy. The experimental results are shown in Figure 3(a). ICN-DIAL-RM significantly outperforms the baseline methods by achieving an average score 2.17 times that of the baselines. One reason for its effectiveness is that, in each round, each of the $N$ agents can engage in $N$ broadcast communications through implicit channels, which significantly enhances information efficiency compared to relying solely on environmental feedback. This result further validates the effectiveness of ICP in more complicated grid world tasks.
@@ 6.3 Hanabi
Because the game rule of Hanabi is quite involved, we refer it to a popular website https://hanabi.github.io/ that discusses both rules and human-play strategies. The action space for each agent includes three options: playing a card out of their 5-card hand, discarding a card, or giving a hint towards another player’s 5-card hand. The hint could reveal all cards in a player’s hand of a specific color or rank. Since they cannot see their own cards, agents must rely on hints from others to infer what cards they hold. The success of the game heavily depends on the agents’ capacity to use these hints effectively to coordinate their actions and play the correct cards in order.
In our experiments, we focus on the 4-player full version of Hanabi. Since hint actions in Hanabi are natural scouting actions, we select them to serve as information carriers for our ICP implementation. We implement several baseline algorithms, including VDN-off-policy and the original version of SAD (Hu & Foerster, 2019), as well as a variant of SAD with the hint action masks removed. For the ICP algorithm, we explore several variations, including ICN-DIAL-RM-on-policy, which involves strategy training with a random initial map using DIAL’s communication gradient technique resulting in an on-policy learning method. We also implement ICN-DIAL-RM-cut-gradient-off-policy, similar to ICN-DIAL-RM-on-policy but with the communication gradient chain cut off to enable off-policy learning. Additionally, we use ICN-RGMComm-DM, which is strategy training with a delayed map using RGMComm (Chen et al., 2024), an off-policy MARL communication algorithm, and ICN-RGMComm-DM-hat-mapping, an extension of ICN-RGMComm-DM that utilizes the hat mapping method for improved communication.
Our presented results are based on the best runs selected from more than three random seeds. From the results shown in Figure 3(b), we observe that the ICP method using the hat mapping approach, ICN-RGMComm-DM-hat-mapping, achieved the best performance. The score achieved by ICP is 24.91, which surpasses the best available learning algorithm (SAD, 23.81 points) and vastly outperforms theory of mind-based methods (Fuchs et al. (2021) 19.13 points) and average human players (17 points, reported in Kantack (2021)). It is worth remarking that the state-of-the-art strategy in the Hanabi game is a search-based algorithm, WTFWThat+search, which reports an averaged score of $24.96$ points (Lerer et al., 2020). With the maximum score of the game being $25$, and there are certain deals that prevent the players from winning $25$ points in any way, both ICP and WTFWThat+search are close to fully solving the game.
For additional details on our implementation, please refer to Appendix A. A comparison between ICP and Hanabi human conventions is provided in Appendix B, along with a detailed analysis of the Hanabi results in Appendix C.
7 Discussions
@@ 7.1 Compatibility of ICP with Other Training Methods
In the ICP framework, action policy allows for the use of various algorithms for training. For instance, the action policy can be trained using value-based methods like VDN (Sunehag et al., 2018) or policy-based methods such as MADDPG (Lowe et al., 2017). This compatibility ensures that ICP can be tailored to various problem settings and agent dynamics. One could choose the most suitable algorithm depending on the specific characteristics of the environment and the agents.
The communication strategy within the ICP framework is also versatile. It can be implemented using a variety of direct channel communication algorithms that support discrete channels. These include well-established methods like DIAL (Foerster et al., 2016) and RGMComm (Chen et al., 2024), which are designed to optimize communication under different constraints and requirements. The flexibility in choosing both the action policy and the communication algorithm highlights the comparability and broad applicability of the ICP framework, making it a general approach for multi-agent reinforcement learning scenarios.
@@ 7.2 Potential of Further Utilizing the Environment Information
As described in Section 4, scouting actions carry two types of information. The ICP framework utilizes the information reflected through the choice of the scouting action (choice information) to establish an implicit communication channel. However, even when we leverage this information, the information reflected through the environment (environment information) remains useful. Although in most cases the constructed implicit channel transmits more stable and useful information, in some situations the environment information is also significant. For example, in the Revealing Goals environment, a revealing action intended to transmit embedded information might also reveal other agents’ goals, which helps agents to make better decisions.
To verify the impact of this environmental information on performance, we randomly shuffle 6 different embeddings on the ICN-RGMComm-hat-mapping policy saved from the previous Hanabi experiment. We then fine-tune the action policy while keeping the information strategy fixed. The results are shown in Figure 4 in the appendix, where, under a fixed information strategy, some randomly shuffled embeddings enjoy improved performance. This suggests that one could further utilize the environment information for more effective communication in future works.
@@ 8 Conclusion and Future Work
The code and our designed environments are freely available in the supplementary material. The details of the implementation can be found in Appendix A. The hyper-parameters used in experiments are also mentioned in Section 6.
Acknowledgement
Baoxiang Wang and Han Wang are partially supported by the National Natural Science Foundation of China (62106213, 72394361), an extended support project from the Shenzhen Science and Technology Program, and support from ByteDance.
@@ A.2 ICP implementation with RGMComm and VDN
A.3 Implementation of DIAL-Cheat and ICN-DIAL-DM
DIAL with a Direct Communication Channel (DIAL-Cheat): In this approach, we added a discrete direct channel to the original environment, with the message size matching the size of the scouting action space. This channel allows agents to broadcast information while simultaneously taking actions.
ICP Implemented with the Delayed Map Approach (ICN-DIAL-DM): For this method, we first removed scouting actions from the environment and introduced a discrete direct channel (with the same message size as the scouting action space). Using this modified environment, we pre-trained the agents with DIAL to learn the information strategy. The learned information strategy was then transferred back to the original environment for fine-tuning, where both the information strategy and action policy were optimized.
@@ Appendix B Human Convention in Hanabi
Compared with H-Group conventions, our ICP framework differs in that the communication strategy itself is learned as part of the training process, rather than being predefined. The communication through informative actions becomes akin to following a learned convention where agents develop their own patterns for encoding and transmitting information during the training phase. This allows agents to use scouting or informative actions as explicit communication mechanisms, while regular actions are selected through an online decision-making policy during gameplay. Notably, our method achieves a score of approximately $24.91$ in the Hanabi game, which significantly surpasses the average human score of around $17$ (Kantack, 2021). Thus, ICP offers a dynamic communication framework where the convention of communication emerges from the agents’ learning process, enabling more flexible and adaptive strategies depending on the environment and task at hand.
Appendix C Analysis of Hanabi results
In Figure 3(b), we present a comparison between our proposed method and SAD, focusing on differences in average scores. While the difference in average score appears minimal in the figure, a detailed analysis provided below reveals that our method achieves a substantial improvement in overall performance.
According to the results reported in the original SAD paper, SAD achieved an average score of 23.81/25 with a win rate of 41.45%. In contrast, our proposed method achieved an average score of 24.91/25 with a win rate of 91%. This represents a more than twofold increase in the win rate. The relatively small difference in average scores, despite the significant improvement in win rates, can be attributed to the score ceiling of 25 in Hanabi.
Hanabi imposes inherent constraints on scoring due to its game structure. Certain starting hands, such as those with multiple high-value cards (e.g., 4s and 5s), can make achieving the maximum score of 25 impossible. Considering these limitations, achieving a win rate of 91%—and scores close to 24 in the remaining 9% of games—demonstrates that our method performs near the theoretical maximum.
Figure 4: By shuffling the embedding of information into scouting actions, even if the information strategy stays the same, environment information will also change. After fine-tuning, the performance of implementation with the same information strategy but different embedding varies.